A brick patio is twice as long as it is wide. It is bordered on all sides by a garden 1.5 m wide. Find the demensions of the partio iff the area of the garden is 54 m
Answer by scott8148(6628)

let x=width, so 2x=length ... (total area)-(patio area)=(garden area)

(x+3)(2x+3)-2x^2=54 ... 2x^2+3x+6x+9-2x^2=54 ... 9x+9=54